OpenStudy (anonymous):

The expression 4n – 1 is a rule to describe a certain sequence. What are the first three terms of the sequence? three, seven, eleven four, eight, twelve three, six, nine four, seven, eleven

OpenStudy (mary.rojas):

check this out http://www.mathsisfun.com/algebra/sequences-finding-rule.html. I dont know how to do this. but i am going to try to help you, just give me a minute

OpenStudy (anonymous):

please hep me

OpenStudy (anonymous):

help

OpenStudy (anonymous):

\(4n-1\) for the first term put \(n=1\) and get \[4\times 1-1=4-1=3\]

OpenStudy (anonymous):

Basically, to create this sequence, you'll take the rule and plug in n=1, then n=2, then n=3, and so on. We only want the first three though.

OpenStudy (anonymous):

for the second term, put \(n=2\) and get \[4\times 2-1=8-1=7\]

OpenStudy (anonymous):

and for the third term put \(n=3\) and see what you get